How do you answer a problem like this?
Twenty-seven exceeds one tenth of a number by seven. What is the number?

Answer by oberobic(2304)

Well, with algebra we always start with 'x'.
x = some number, we just don't know what number yet
.
1/10*x +7 = 27
1/10*x = 20
x = 200
.
Hmmm. We're still not sure.
so...
.
We ALWAYS check our work!
.
1/10*200 = 20
27 - 20 = 7
So, "27 exceeds one tenth of a number by 7."
Correct.
.
Answer: The number is 200.
.
Done.